OpenStudy (anonymous):

John has 900 dollars in his savings account for extra expenses. He is spending 75 dollars each month. Let x= months. Part A. Write a function to show Johns savings account at the end of the first month. Part B. How much money does John have left after 10 months? Part C. How many months until John has no money left in his account? Help please!!!!

OpenStudy (anonymous):

900-75x would be the overall equation A. 900-75(1) 900-75=825 B. 900-75(10) 900-750=150 C. 900-75x=0 -75x=-900 x=12 In one year John will be out of money

OpenStudy (anonymous):

Wow That was quick. Thanks!